OpenStudy (dtan5457):

finding all 6 trig functions question

OpenStudy (dtan5457):

if given cot theta=6/11, knowing that theta is a acute angle. could sin be both

OpenStudy (jhannybean):

You could always change everything in terms of sines and cosines.

OpenStudy (dtan5457):

\[\frac{ 6\sqrt{157} }{ 157 }\] and \[\frac{ 11\sqrt{157} }{ 157 }\]

OpenStudy (dtan5457):

depending on where you set up the triangle

OpenStudy (dtan5457):

yes?

OpenStudy (rational):

Also draw a triangle

OpenStudy (dtan5457):